# Solvable Tuple Patterns and Their Applications to Program Verification

### v1

@@ 2.1 Tuple Patterns and Inference Algorithm
When $\Theta t=M$ holds, we also write $M,\Theta\models t$. We also write $M,\Theta\models_{s}t$ if $\Theta t=M$ and there exists no variable $x$ such that $\Theta(x)=\widetilde{\epsilon}$ (i.e., if no variable is always mapped to $\epsilon$). We write $M\models t$ ($M\models_{s}t$, resp.) if there exists $\Theta$ such that $M,\Theta\models t$ ($M,\Theta\models_{s}t$, resp.), and we call $\Theta$ a witness substitution for $M\models t$ ($M\models_{s}t$, resp.).
@@
Let $M$ be: $\left(\begin{array}[]{ll}a&aa\\ b&bb\end{array}\right)$ and $t$ be $(x,xx)$. Then $M\models_{s}t$, where $\Theta=\left(\begin{array}[]{c}x\\ a\\ b\end{array}\right)$ is a witness substitution. We also have $M,\Theta\models(x,xxy)$ for $\Theta=\left(\begin{array}[]{cc}x&y\\ a&\epsilon\\ b&\epsilon\end{array}\right)$, but $M\not\models_{s}(x,xxy)$ as $\Theta$ always maps $y$ to $\epsilon$. ∎
@@
The tuple pattern inference problem is the problem of, given learning data $M$, finding a tuple pattern $(p_{1},\ldots,p_{n})$ such that $M\models(p_{1},\ldots,p_{n})$. For instance, for $M$ in Example 2.2, $(ax,y,axy)$ is an answer to the inference problem.
2.1.3 Inference Algorithm.
We now present a (non-deterministic) algorithm for tuple pattern inference. We formalize it by using a rewriting relation $(t,\Theta)\longrightarrow(t^{\prime},\Theta^{\prime})$. Starting with the initial state $((x_{1},\ldots,x_{n}),[M/(x_{1},\ldots,x_{n})])$, the algorithm repeatedly applies the rewriting rules below until no further rule is applicable. When the final pair is $(t^{\prime},\Theta^{\prime})$, then $t^{\prime}$ is the inferred tuple pattern. In fact, as we will see later (in Theorem 2.1), $((x_{1},\ldots,x_{n}),[M/(x_{1},\ldots,x_{n})])\longrightarrow^{*}(t^{\prime},\Theta^{\prime})$ implies $M,\Theta^{\prime}\models t^{\prime}$, because the rewriting relation $(t,\Theta)\longrightarrow(t^{\prime},\Theta^{\prime})$ maintains the invariant $\Theta t=\Theta^{\prime}t^{\prime}$.
$\frac{\begin{array}[]{@{}c@{}}M[*][i]\neq\widetilde{\epsilon}\quad\quad M[*][j]=M[*][i]\cdot\widetilde{s}\quad\quad x_{j}^{\prime}\mbox{ fresh}\end{array}}{\begin{array}[]{@{}c@{}}(t,[M/(x_{1},\ldots,x_{m})])\longrightarrow([x_{i}\cdot x_{j}^{\prime}/x_{j}]t,[M\{j\mapsto\widetilde{s}\}/(x_{1},\ldots,x_{j-1},x_{j}^{\prime},x_{j+1},\ldots,x_{m})])\end{array}}$ (R-Prefix)
@@
The rule R-Prefix is for the case where the $i$-th column of $M$ is a prefix of the $j$-th column. In this case, we replace the $j$-th column with its postfix obtained by removing the prefix M[][i]M[*][i], and instantiate $x_{j}$ in $t$ accordingly to $x_{i}\cdot x_{j}^{\prime}$. The rule R-CPrefix is for the case where a constant $a$ is a prefix of the $j$-th column, and the rule R-Epsilon is for the case where every element of the $j$-th column is an empty sequence. In the latter case, the $j$-th column is removed.
@@
$((x_{1},x_{2},x_{3}),[M/(x_{1},x_{2},x_{3})])\longrightarrow_{\mbox{(\hbox{\sc R-Prefix}, $x_{3}=x_{1}x_{3}^{\prime}$)}}((x_{1},x_{2},x_{1}x_{3}^{\prime}),\left(\begin{array}[]{rrr}x_{1}&x_{2}&x_{3}^{\prime}\\ a&\ b&\ b\\ aa&\epsilon&\epsilon\end{array}\right))$
 $\longrightarrow_{\mbox{ (\hbox{\sc R-Prefix}, $x_{3}^{\prime}=x_{2}x_{3}^{\prime\prime}$)}}((x_{1},x_{2},x_{1}x_{2}x_{3}^{\prime\prime}),\left(\begin{array}[]{rrr}x_{1}&x_{2}&x_{3}^{\prime\prime}\\ a&\ b&\ \epsilon\\ aa&\epsilon&\epsilon\end{array}\right))$
 $\longrightarrow_{\mbox{ (\hbox{\sc R-Epsilon}, $x_{3}^{\prime\prime}=\widetilde{\epsilon}$)}}((x_{1},x_{2},x_{1}x_{2}),\left(\begin{array}[]{rr}x_{1}&x_{2}\\ a&\ b\\ aa&\epsilon\end{array}\right))$
 $\longrightarrow_{\mbox{ (\hbox{\sc R-CPrefix}, $x_{1}=ax_{1}^{\prime}$)}}((ax^{\prime}_{1},x_{2},ax^{\prime}_{1}x_{2}),\left(\begin{array}[]{rr}x_{1}^{\prime}&x_{2}\\ \epsilon&\ b\\ a&\epsilon\end{array}\right))\;\mathbin{\ \not\!\!\longrightarrow}{}$
Actually, the patterns such as $(x_{1},x_{2},x_{1}x_{3}^{\prime})$ and $(x_{1},x_{2},x_{1}x_{2})$ obtained in intermediate steps are also valid solutions of the inference problem, but the final pattern describes a smaller relation than them. ∎
@@
$((x_{1},x_{2},x_{3}),[M/(x_{1},x_{2},x_{3})])\longrightarrow((x_{1},x_{2},x_{1}x_{3}^{\prime}),\left(\begin{array}[]{rrr}x_{1}&x_{2}&x_{3}^{\prime}\\ aa&\ a&\ c\\ b&\ bb&\ bd.\end{array}\right))$
$((x_{1},x_{2},x_{3}),[M/(x_{1},x_{2},x_{3})])\longrightarrow((x_{1},x_{2},x_{2}x_{3}^{\prime}),\left(\begin{array}[]{rrr}x_{1}&x_{2}&x_{3}^{\prime}\\ aa&\ a&\ ac\\ b&\ bb&\ d.\end{array}\right)).$
The first pattern obtained describes that the first element is a prefix of the third element, while the second patter describes that the second element is a prefix of the third element. Unfortunately, there is no single tuple pattern that describes both properties; we need to introduce a conjunctive tuple pattern like $(x,y,xz)\land(u,v,vw)$, which is out of the scope of this paper. ∎
We now state some key properties of the algorithm; other important properties are discussed in Section 2.2. Proofs omitted below are found in Appendix A. The following theorem states the soundness of the algorithm.
@@
The theorem follows immediately from the fact that $(t_{1},\Theta_{1})\longrightarrow(t_{2},\Theta_{2})$ implies $\Theta_{1}t_{1}=\Theta_{2}t_{2}$. See Appendix A for more details.
@@
Given $M$ as input, the algorithm above outputs $t$ such that $(\widetilde{x},[M/\widetilde{x}])\longrightarrow^{*}(t,\Theta)\;\mathbin{\ \not\!\!\longrightarrow}{}$ in time polynomial in $\mathit{size}(M)$.
@@
2.2 Solvable Tuple Patterns
Not all the tuple patterns are inferable by the algorithm described above. For example, the pattern $(xx)$ on a singleton tuple is obviously non-inferable. In this subsection, we introduce the notion of solvable tuple patters (STPs), to characterize the class of tuple patterns that can be inferred by our algorithm.
@@ 2.2 Solvable Tuple Patterns
Instead of restricting the class of tuple patterns, one may expect to obtain an extension of the algorithm that can infer arbitrary tuple patterns. Unfortunately, learning arbitrary tuple patterns from positive samples is computationally infeasible. As mentioned in Remark 2.1, singleton tuple patterns coincide with E-patterns [49], and the full class of E-pattern languages is not learnable from positive data in general (more precisely, it is not learnable when the alphabet size is 2, 3, or 4 [44, 45] and the learnability is unknown when the alphabet size is larger). Also, the membership problem (of deciding whether a string $s$ matches pattern $p$) is NP-complete for both the original pattern languages (NE-pattern languages) [1] and E-pattern languages [27, 42], and the inclusion problem for E-pattern languages is undecidable [26, 42]. ∎
Let $\mathcal{T}_{k}$ be the set:
$\{t\mid t\leadsto^{*}(x_{1},\ldots,x_{n}),|t|=k,\mbox{ and $x_{1},\ldots,x_{n}$ are distinct from each other}\},$
where $t_{1}\leadsto t_{2}$ is defined by:
$\frac{\begin{array}[]{@{}c@{}}p_{j}=p_{i}\cdot p^{\prime}_{j}\quad\quad p_{i}\neq\epsilon\end{array}}{\begin{array}[]{@{}c@{}}(p_{1},\ldots,p_{n})\leadsto(p_{1},\ldots,p_{j-1},p^{\prime}_{j},p_{j+1},\ldots,p_{n})\end{array}}$ (PR-Prefix)
$\frac{\begin{array}[]{@{}c@{}}p_{j}=a\cdot p^{\prime}_{j}\quad\quad a\in\Sigma\end{array}}{\begin{array}[]{@{}c@{}}(p_{1},\ldots,p_{n})\leadsto(p_{1},\ldots,p_{j-1},p^{\prime}_{j},p_{j+1},\ldots,p_{n})\end{array}}$ (PR-CPrefix)
$\frac{\begin{array}[]{@{}c@{}}p_{j}=\epsilon\end{array}}{\begin{array}[]{@{}c@{}}(p_{1},\ldots,p_{n})\leadsto(p_{1},\ldots,p_{j-1},p_{j+1},\ldots,p_{n})\end{array}}$ (PR-Epsilon)
We write $\mathcal{T}$ for $\bigcup_{k}\mathcal{T}_{k}$. A tuple pattern $t$ of arity $k$ is called solvable if $t\in\mathcal{T}_{k}$.
Each rule of name PR-xx corresponds to the rule R-xx; while the rule R-xx manipulates data $M$, the rule PR-xx applies the corresponding operation on a tuple pattern, and simplifies the pattern. A tuple pattern $t$ belongs to $\mathcal{T}$ if $t$ can be reduced to a trivial pattern $(x_{1},\ldots,x_{n})$. We call the pattern modified or removed in the rewriting $(p_{1},\ldots,p_{n})\leadsto(p^{\prime}_{1},\ldots,p^{\prime}_{m})$ (i.e., $p_{j}$ in the above rules) a principal pattern. We also call the pattern $p_{i}$ in PR-Prefix an auxiliary pattern.

### v2

@@ 2.2 Tuple Pattern Inference Problem
When $\Theta t=M$ holds, we also write $M,\Theta\models t$. We also write $M,\Theta\models_{s}t$ if $\Theta t=M$ and there exists no variable $x$ such that $\Theta(x)=\widetilde{\epsilon}$ (i.e., if no variable is always mapped to $\epsilon$). We write $M\models t$ ($M\models_{s}t$, resp.) if there exists $\Theta$ such that $M,\Theta\models t$ ($M,\Theta\models_{s}t$, resp.), and we call $\Theta$ a witness substitution for $M\models t$ ($M\models_{s}t$, resp.). For a conjunctive tuple pattern $\gamma=t_{1}\land\cdots\land t_{m}$, we write $M\models\gamma$ ($M\models_{s}\gamma$, resp.) if $M\models t_{i}$ ($M\models_{s}t_{i}$, resp.) holds for every $i\in\{1,\ldots,m\}$.
@@
Let $M$ be: $\left(\begin{array}[]{ll}a&aaa\\ b&abb\end{array}\right)$ and $t$ be $(x,axx)$. Then $M\models_{s}t$, where $\Theta=\left(\begin{array}[]{c}x\\ a\\ b\end{array}\right)$ is a witness substitution. We also have $M,\Theta\models(x,axxy)$ for $\Theta=\left(\begin{array}[]{cc}x&y\\ a&\epsilon\\ b&\epsilon\end{array}\right)$, but $M\not\models_{s}(x,xxy)$ as $\Theta$ always maps $y$ to $\epsilon$. ∎
@@
The tuple pattern inference problem is the problem of, given learning data $M$, finding a tuple pattern $(p_{1},\ldots,p_{n})$ such that $M\models(p_{1},\ldots,p_{n})$. For instance, for $M$ in Example 2.1, $(x,axx)$ is an answer to the inference problem.
2.3 Inference Algorithm
[…]
$\frac{\begin{array}[]{@{}c@{}}M[*][i]\neq\widetilde{\epsilon}\quad\quad M[*][j]=M[*][i]\cdot\widetilde{s}\quad\quad x_{j}^{\prime}\mbox{ fresh}\end{array}}{\begin{array}[]{@{}c@{}}(t,[M/(x_{1},\ldots,x_{m})])\longrightarrow([x_{i}x_{j}^{\prime}/x_{j}]t,[M\{j\mapsto\widetilde{s}\}/(x_{1},\ldots,x_{j-1},x_{j}^{\prime},x_{j+1},\ldots,x_{m})])\end{array}}$ (R-Prefix)
@@ 2.3 Inference Algorithm
$\frac{\begin{array}[]{@{}c@{}}M[*][i]\neq\widetilde{\epsilon}\quad\quad M[*][j]=\widetilde{s}\cdot M[*][i]\quad\quad x_{j}^{\prime}\mbox{ fresh}\end{array}}{\begin{array}[]{@{}c@{}}(t,[M/(x_{1},\ldots,x_{m})])\longrightarrow([x_{j}^{\prime}x_{i}/x_{j}]t,[M\{j\mapsto\widetilde{s}\}/(x_{1},\ldots,x_{j-1},x_{j}^{\prime},x_{j+1},\ldots,x_{m})])\end{array}}$ (R-Suffix)
$\frac{\begin{array}[]{@{}c@{}}M[*][j]=\widetilde{s}\cdot a\quad\quad a\in\Sigma\quad\quad x_{j}^{\prime}\mbox{ fresh}\end{array}}{\begin{array}[]{@{}c@{}}(t,[M/(x_{1},\ldots,x_{m})])\longrightarrow([x_{j}^{\prime}a/x_{j}]t,[M\{j\mapsto\widetilde{s}\}/(x_{1},\ldots,x_{j-1},x_{j}^{\prime},x_{j+1},\ldots,x_{m})])\end{array}}$ (R-CSuffix)
@@
The rule R-Prefix is for the case where the $i$-th column of $M$ is a prefix of the $j$-th column. In this case, we replace the $j$-th column with its suffix obtained by removing the prefix M[][i]M[*][i], and instantiate $x_{j}$ in $t$ accordingly to $x_{i}\cdot x_{j}^{\prime}$. The rule R-CPrefix is for the case where a constant $a$ is a prefix of the $j$-th column, and the rule R-Epsilon is for the case where every element of the $j$-th column is an empty sequence. In the latter case, the $j$-th column is removed.
Example 2.3.
Recall the data in Example 2.1: $M=\left(\begin{array}[]{ll}a&aaa\\ b&abb\end{array}\right)$. The pattern $(x,axx)$ is inferred (up to variable renaming) as follows.
$\Big((x_{1},x_{2}),\left(\begin{array}[]{rr}x_{1}&x_{2}\\ a&\ aaa\\ b&abb\end{array}\right)\Big)\longrightarrow\Big((x_{1},ax_{2}^{\prime}),\left(\begin{array}[]{rr}x_{1}&x_{2}^{\prime}\\ a&\ aa\\ b&bb\end{array}\right)\Big)\longrightarrow\Big((x_{1},ax_{1}x_{2}^{\prime\prime}),\left(\begin{array}[]{rr}x_{1}&x_{2}^{\prime\prime}\\ a&\ a\\ b&\ b\end{array}\right)\Big)$
$\longrightarrow\Big((x_{1},ax_{1}x_{1}x_{2}^{\prime\prime\prime}),\left(\begin{array}[]{rr}x_{1}&x_{2}^{\prime\prime\prime}\\ a&\ \epsilon\\ b&\epsilon\end{array}\right)\Big)\longrightarrow\Big((x_{1},ax_{1}x_{1}),\left(\begin{array}[]{cc}x_{1}\\ a\\ b\end{array}\right)\Big)\;\mathbin{\ \not\!\!\longrightarrow}{}$
Actually, the patterns such as $(x_{1},ax_{2}^{\prime})$ and $(x_{1},ax_{1}x_{1}x_{2}^{\prime\prime\prime})$ obtained in intermediate steps are also valid solutions of the inference problem, but the final pattern describes a smaller relation than them. ∎
@@
$\Big((x_{1},x_{2},x_{3}),\left(\begin{array}[]{rrr}x_{1}&x_{2}&x_{3}\\ a&\ b&\ ab\\ aa&\epsilon&aa\end{array}\right)\Big)\longrightarrow_{\mbox{(\hbox{\sc R-Prefix}, $x_{3}=x_{1}x_{3}^{\prime}$)}}\Big((x_{1},x_{2},x_{1}x_{3}^{\prime}),\left(\begin{array}[]{rrr}x_{1}&x_{2}&x_{3}^{\prime}\\ a&\ b&\ b\\ aa&\epsilon&\epsilon\end{array}\right)\Big)$
 $\longrightarrow_{\mbox{ (\hbox{\sc R-Prefix}, $x_{3}^{\prime}=x_{2}x_{3}^{\prime\prime}$)}}\Big((x_{1},x_{2},x_{1}x_{2}x_{3}^{\prime\prime}),\left(\begin{array}[]{rrr}x_{1}&x_{2}&x_{3}^{\prime\prime}\\ a&\ b&\ \epsilon\\ aa&\epsilon&\epsilon\end{array}\right)\Big)$
 $\longrightarrow_{\mbox{ (\hbox{\sc R-Epsilon}, $x_{3}^{\prime\prime}=\widetilde{\epsilon}$)}}\Big((x_{1},x_{2},x_{1}x_{2}),\left(\begin{array}[]{rr}x_{1}&x_{2}\\ a&\ b\\ aa&\epsilon\end{array}\right)\Big)$
 $\longrightarrow_{\mbox{ (\hbox{\sc R-CPrefix}, $x_{1}=ax_{1}^{\prime}$)}}\Big((ax^{\prime}_{1},x_{2},ax^{\prime}_{1}x_{2}),\left(\begin{array}[]{rr}x_{1}^{\prime}&x_{2}\\ \epsilon&\ b\\ a&\epsilon\end{array}\right)\Big)\;\mathbin{\ \not\!\!\longrightarrow}{}$
@@
$\Big((x_{1},x_{2},x_{3}),\left(\begin{array}[]{rrr}x_{1}&x_{2}&x_{3}\\ aa&\ a&\ aac\\ b&\ bb&\ bbd.\end{array}\right)\Big)\longrightarrow\Big((x_{1},x_{2},x_{1}x_{3}^{\prime}),\left(\begin{array}[]{rrr}x_{1}&x_{2}&x_{3}^{\prime}\\ aa&\ a&\ c\\ b&\ bb&\ bd.\end{array}\right)\Big)$
$\Big((x_{1},x_{2},x_{3}),\left(\begin{array}[]{rrr}x_{1}&x_{2}&x_{3}\\ aa&\ a&\ aac\\ b&\ bb&\ bbd.\end{array}\right)\Big)\longrightarrow\Big((x_{1},x_{2},x_{2}x_{3}^{\prime}),\left(\begin{array}[]{rrr}x_{1}&x_{2}&x_{3}^{\prime}\\ aa&\ a&\ ac\\ b&\ bb&\ d.\end{array}\right)\Big).$
The first pattern obtained describes that the first element is a prefix of the third element, while the second patter describes that the second element is a prefix of the third element. The conjunctive pattern $(x_{1},x_{2},x_{1}x_{3}^{\prime})\land(x_{1},x_{2},x_{2}x_{3}^{\prime})$ subsumes both patterns. ∎
We now state some key properties of the algorithm; other important properties are discussed in Section 3. Proofs omitted below are found in Appendix A. The theorem below states the soundness of the algorithm, which follows immediately from the fact that $(t_{1},\Theta_{1})\longrightarrow(t_{2},\Theta_{2})$ implies $\Theta_{1}t_{1}=\Theta_{2}t_{2}$. See Appendix A for more details.
@@
Given $M$ as input, $\mathtt{TPinf}$ outputs $t$ such that $(\widetilde{x},[M/\widetilde{x}])\longrightarrow^{*}(t,\Theta)\;\mathbin{\ \not\!\!\longrightarrow}{}$ in time polynomial in $\mathit{size}(M)$.
@@
An algorithm $\mathtt{CTPinf}$ to infer a conjunctive tuple pattern is given by: $\mathtt{CTPinf}(M)=\bigwedge\{t^{\prime}\mid(\widetilde{x},[M/\widetilde{x}])\longrightarrow^{*}(t,\Theta)\;\mathbin{\ \not\!\!\longrightarrow}\}$. Note that the set of $(t,\Theta)$ such that $(\widetilde{x},[M/\widetilde{x}])\longrightarrow^{*}(t,\Theta)$ is finite; thus we can indeed compute $\mathtt{CTPinf}(M)$ by an exhaustive search. $\mathtt{CTPinf}$, however, runs in time exponential in $\mathit{size}(M)$ in the worst case.
3 Solvable Tuple Patterns
In this section, we introduce the notions of solvable tuple patterns (STPs) and conjunctive solvable tuple patters (CSTPs), which characterize the classes of (conjunctive) tuple patterns that can be inferred by our algorithms. The pattern $(xx)$ on a singleton tuple is obviously non-inferable by our algorithm, and it is indeed deemed non-solvable in the characterization below. In contrast, the pattern $(xx,x)$ is solvable.
@@ 3 Solvable Tuple Patterns
Instead of restricting the class of tuple patterns, one may expect to obtain an extension of the algorithm that can infer arbitrary tuple patterns. Unfortunately, learning arbitrary tuple patterns from positive samples is computationally infeasible. As mentioned in Remark 2.1, singleton tuple patterns coincide with E-patterns [57], and the full class of E-pattern languages is not learnable from positive data in general (more precisely, it is not learnable when the alphabet size is 2, 3, or 4 [52, 53] and the learnability is unknown when the alphabet size is larger). Also, the inclusion problem is undecidable for both NE- and E-pattern languages [28, 47]. In contrast, STPs have good algorithmic properties, as discussed in Section 3.4. ∎
3.1 Solvability
Let $\mathcal{T}_{k}$ be the set: $\{t\mid t\leadsto^{*}(x_{1},\ldots,x_{n}),|t|=k,\mbox{ and $x_{1},\ldots,x_{n}$ are distinct from each other}\}$, where $t_{1}\leadsto t_{2}$ is defined by:
$\frac{\begin{array}[]{@{}c@{}}p_{j}=p_{i}\cdot p^{\prime}_{j}\quad\quad p_{i}\neq\epsilon\end{array}}{\begin{array}[]{@{}c@{}}(p_{1},\ldots,p_{n})\leadsto(p_{1},\ldots,p_{j-1},p^{\prime}_{j},p_{j+1},\ldots,p_{n})\end{array}}$ (PR-Prefix) $\frac{\begin{array}[]{@{}c@{}}p_{j}=a\cdot p^{\prime}_{j}\quad\quad a\in\Sigma\end{array}}{\begin{array}[]{@{}c@{}}(p_{1},\ldots,p_{n})\leadsto(p_{1},\ldots,p_{j-1},p^{\prime}_{j},p_{j+1},\ldots,p_{n})\end{array}}$ (PR-CPrefix) $\frac{\begin{array}[]{@{}c@{}}p_{j}=p^{\prime}_{j}\cdot p_{i}\quad\quad p_{i}\neq\epsilon\end{array}}{\begin{array}[]{@{}c@{}}(p_{1},\ldots,p_{n})\leadsto(p_{1},\ldots,p_{j-1},p^{\prime}_{j},p_{j+1},\ldots,p_{n})\end{array}}$ (PR-Suffix) $\frac{\begin{array}[]{@{}c@{}}p_{j}=p^{\prime}_{j}\cdot a\quad\quad a\in\Sigma\end{array}}{\begin{array}[]{@{}c@{}}(p_{1},\ldots,p_{n})\leadsto(p_{1},\ldots,p_{j-1},p^{\prime}_{j},p_{j+1},\ldots,p_{n})\end{array}}$ (PR-CSuffix) $\frac{\begin{array}[]{@{}c@{}}p_{j}=\epsilon\end{array}}{\begin{array}[]{@{}c@{}}(p_{1},\ldots,p_{n})\leadsto(p_{1},\ldots,p_{j-1},p_{j+1},\ldots,p_{n})\end{array}}$ (PR-Epsilon)
We write $\mathcal{T}$ for $\bigcup_{k}\mathcal{T}_{k}$. A tuple pattern $t$ of arity $k$ is called solvable if $t\in\mathcal{T}_{k}$. A conjunctive tuple pattern $\gamma=t_{1}\land\cdots\land t_{m}$ is solvable if $t_{1},\ldots,t_{m}\in\mathcal{T}_{k}$ for some $k$. Each rule of name PR-xx corresponds to the rule R-xx; while the rule R-xx manipulates data $M$, the rule PR-xx applies the corresponding operation to a tuple pattern, and simplifies the pattern. A tuple pattern $t$ belongs to $\mathcal{T}$ if $t$ can be reduced to a trivial pattern $(x_{1},\ldots,x_{n})$. We call the pattern modified or removed in the rewriting $(p_{1},\ldots,p_{n})\leadsto(p^{\prime}_{1},\ldots,p^{\prime}_{m})$ (i.e., $p_{j}$ in the above rules) a principal pattern. We also call the pattern $p_{i}$ in PR-Prefix or PR-Suffix an auxiliary pattern.